I am trying to solve a linear system by elimination but am coming up with a problem. At least I think it is a problem.
2x-5y=11
3x+y=8
I will multiply by (3) and (-2) to get 6x and -6x
(3)2x-5y=11
(-2)3x+y=8
6x-5y=11
-6x-y=8
subtract 5y+1y=6y
11 + 8 =19
6y=19
When I divide 19 by 6 I get 3.16. What did I do wrong????? Please help.
Thank you so much. Tricia

Answer by indiangangsta(5)

your problem is a very common one. i will show the steps below:
2x-5y=11
3x+y=8
multiply by 3 and -2
(3)2x-5y=11
(-2)3x+y=8
Now look carefully, and u'll find ur mistake:
6x-15y=33 --> you have to multiply the 2, the -5, and the 11 by 3, which u forgot
-6x+-2y=-16 --> likewise, u have to multiply the 3, 1, and 8 by -2
from there, you just continue on the regular course of action, and u'll get the right answer.